In a certain men's track and field event, the shotput has a mass of 7.25 kg and is released eastward with a speed of 17.2 m/s at 41.0
Answer #1

The initial horizontal components of the momentum of this shotput is,

                 Px = mvcos41o = 7.25*17.2*cos41o = 94.1 kg.m/s

The initial vertical components of the momentum of this shotput is,

                 Py = mvsin41o = 7.25*17.2*sin41o = 81.8 kg.m/s

The initial momentum is directed at 41o above the horizontal.